# On manifolds with almost non-negative Ricci curvature and integrally-positive $k^{th}$-scalar curvature

### v2

@@ 2 Preliminaries
The next proposition, from [KL, Theorem $1.1$], characterizes the $\mathsf{RCD}$ spaces having essential dimension $1$.
@@
We now recall some topological properties of $\mathsf{RCD}$ spaces. The next proposition is a special case of [Santos, Theorem 5], in the non-collapsing scenario. We denote by $\mathrm{b}_{1}(X)$ the first Betti number of $X$.
@@
The next proposition, proved in [RCDsemilocally, Corollary 1.2] (see also [wang2021riccilimitspacessemilocally, Theorem 1.2]), establishes the converse inequality in higher generality.
@@
A connected metric space $(\tilde{X},\mathsf{d}_{\tilde{X}})$ is said to be a universal cover for $(X,\mathsf{d}_{X})$ if it is a covering space with the following property: for any other connected covering space $(Y,\mathsf{d}_{Y})$ of $(X,\mathsf{d}_{X})$, there exists a continuous map $f:\tilde{X}\to Y$ such that the triangle made with the projection maps onto $X$ commutes.
The next result proved in [MondinoWei] establishes the existence of the universal cover for an $\mathsf{RCD}(K,N)$ space.
@@
Let $(X,\mathsf{d},\mathfrak{m})$ be an $\mathsf{RCD}(K,N)$ space. Then, it admits a universal cover $(\tilde{X},\mathsf{d}_{\tilde{X}})$. Moreover there exists a measure $\tilde{\mathfrak{m}}$ on $\tilde{X}$ such that $(\tilde{X},\mathsf{d}_{\tilde{X}},\tilde{\mathfrak{m}})$ is an $\mathsf{RCD}(K,N)$ space.
Since $\mathsf{RCD}(K,N)$ spaces are semi-locally simply connected thanks to [RCDsemilocally] the universal is simply connected. The next proposition follows from [MondinoWei, Theorem 1.3], after taking into account that the revised fundamental group mentioned in the reference coincides with the usual fundamental group thanks to [RCDsemilocally].
@@
The next proposition from [Sormaniwei, Corollary 4.11] shows that the universal cover of a Ricci limit space is a Ricci limit space, as well.
@@
Let $K\in\mathbb{R}$ and $n\in\mathbb{N}$. Let $(M_{j}^{n},g_{j},\bar{x}_{j})$ be a sequence of Riemannian manifolds with $\mathsf{Ric}_{M_{j}}\geq K$ converging in pGH-sense to a compact metric space $(X,\mathsf{d})$. Let $(\tilde{X},\tilde{\mathsf{d}})$ be the universal cover of $X$ and fix $\tilde{x}\in\tilde{X}$. Then, there is a sequence of coverings $(\tilde{M}_{j},\tilde{g}_{j},\tilde{x}_{j})$ of $M_{j}$ converging in pGH-sense to $(\tilde{X},\tilde{\mathsf{d}},\tilde{x})$.
The next proposition follows combining [Anderson, Theorems $1.1$ and $1.3$] with the fact that each covering of a fixed manifold has volume growth controlled from above by the one of the universal cover.
@@
Let $k,h\in\mathbb{N}\cup\{0\}$ with $k\leq h$. Let $(M,g,\bar{x})$ be a pointed Riemannian manifold with $\mathsf{Vol}(B_{t}(\bar{x}))\geq c_{1}t^{k}$ for some $c_{1}>0$ and whose universal cover $(\tilde{M},\tilde{g},\tilde{x})$ satisfies $\widetilde{\mathsf{Vol}}(B_{t}(\tilde{x}))\leq c_{2}t^{h}$ for some $c_{2}>0$. Then $\mathrm{b}_{1}(M)\leq h-k$.
@@ 3 Integral estimates on scalar curvature and first geometric applications
Let $k,n\in\mathbb{N}$ with $k\leq n$, and let $c,\epsilon>0$ be fixed. There exist $c_{1}(n,\epsilon),\epsilon^{\prime}(n,\epsilon)>0$ arbitrarily small satisfying the following. Let $A=(a_{j,i})\in\mathbb{R}^{n\times k}$ be a matrix with $|A^{T}A-I_{k}|\leq c_{1}$ and $|A|^{2}\leq c$. Consider real numbers $-\epsilon^{\prime}\leq\lambda_{1}\leq\cdots\leq\lambda_{n}$ with $\lambda_{k}\geq\epsilon/2k$. Then
@@
Assume by contradiction that the statement fails. Hence, we find sequences $c_{h},\epsilon^{\prime}_{h}\downarrow 0$ and $A_{h}=(a^{h}_{j,i})\in\mathbb{R}^{n\times k}$ with $|A^{T}_{h}A_{h}-I_{k}|\leq c_{h}$, $|A_{h}|^{2}\leq c$ and real numbers $-\epsilon^{\prime}_{h}\leq\lambda^{h}_{1}\leq\cdots\leq\lambda^{h}_{n}$ with $\lambda_{k}\geq\epsilon/2k$, with
@@
Moreover, since $|A|^{2}\leq c$, $-\epsilon^{\prime}_{h}\leq\lambda^{h}_{1}\leq\cdots\leq\lambda^{h}_{n}$ and $\lambda^{h}_{k}\geq\epsilon/2k$, it holds
@@
The last two identities are clearly in contradiction: indeed, the former states that the columns of $A$ are $k$ orthonormal vectors in $\mathbb{R}^{n}$, while the latter implies that their span is at most $k-1$ dimensional. ∎
@@
Observe that the condition in the previous definition is independent of the chosen orthonormal basis for $T_{p}M$. Fix now an orthonormal basis in $T_{p}M$ such that $\mathsf{Ric}_{M}$ in $T_{p}M$ is represented by the diagonal matrix $\Lambda\in\mathbb{R}^{n\times n}$, whose eigenvalues are $-\epsilon^{\prime}\leq\lambda_{1}\leq\cdots\leq\lambda_{n}$.
@@
00 0 $\vee\sum_{j=1}^{k}\mathsf{Ric}(\nabla u_{j},\nabla u_{j})=0\vee\sum_{j=1}^{k}\nabla u_{j}^{T}\Lambda\nabla u_{j}$
@@
We first prove that $E^{\prime}\subset E$. It is sufficient to show that at every point of $E^{\prime}$ the quantity $|\nabla u^{T}\nabla u-I_{k}|$ is smaller than the constant $c_{1}(\epsilon,n)$ appearing in the definition of $E$. We prove it by a telescopic argument (cf. [MonBruSem]). Set $\mathcal{E}:=\sum_{i,j}|\nabla u_{j}\cdot\nabla u_{j}-\delta_{ij}|$. For every $j\in\mathbb{N}$ and $y\in B_{1}(x)$, we have
@@
Choosing $\eta$ and $\epsilon^{\prime}$ to be sufficiently small, and letting $j$ go to +$\infty$+ ∞, we conclude that $E^{\prime}\subset E$.
@@
To prove this, we use a Vitali covering argument coupled with a scale-picking (cf. [MonBruSem]). For every $y\in B_{1}(x)$ such that
@@
[…]
converge to 00 as $\delta\searrow 0$.
@@
Let $k,n\in\mathbb{N}$ with $k<n$. We show that for every $20<\epsilon<1/$ 2 there exists $\delta>0$ such that if $(M^{n},g)$ has $\mathsf{Ric}_{M}\geq-\delta$ and $B_{10}(x)\subset M$ is $(k,\delta)$-symmetric, then
@@
We next show (11). To this aim, fix $s\in(0,1)$ and consider the case $k=n$. We want show that for every $20<\epsilon<1/$ 2 there exists $\delta>0$ such that if $(M^{n},g,x)$ has $\mathsf{Ric}_{M}\geq-\delta$ and $B_{10}(x)\subset M$ is $\delta$-regular, then
@@
[…]
as $\delta\to 0$. This would allow to obtain a priori integral bounds on the scalar curvature outside of the quantitative singular set of a manifold with Ricci curvature bounded from below. Unfortunately, it is shown in [DephilippisHarmonic] that such estimates cannot exist in such generality.
@@ 3.2 First geometric applications in almost non-negative Ricci and integrally-positive scalar curvature
We now obtain the first geometric applications of the results of the previous section. We introduce some notation first (cf. [ShenKricci]).
@@
Let $(M^{n},g)$ be a manifold, let $11\leq k\leq n-$ 1, let $x\in M$, and let $c\in\mathbb{R}$. We denote by $R_{M}(\cdot,\cdot,\cdot,\cdot)$ the Riemann curvature tensor of $M$. We say that $\mathsf{Ric}_{M,k}\geq c$ (resp. $\leq c$) in $T_{x}M$ if, for all $(k+1)$-dimensional subspaces $V\subset T_{x}M$, it holds
@@
Let $(M^{n},g)$ be a manifold, let $11\leq k\leq n-$ 1, let $A\subset M$ be measurable, let $c\in\mathbb{R}$, and let $F:\mathbb{R}\to\mathbb{R}$ be a measurable function. We say that
[…]
@@
[…]
The next theorem should be compared with [wang2023positive, Theorem 1.1], where similar conclusions were obtained under stronger assumptions: non-negative Ricci curvature (below, almost non-negative Ricci curvature) and scalar curvature greater than 2 (below, a local integral lower bound on the scalar curvature). Whenever we consider a sequence of manifolds $(M_{j},g_{j})$, we write $\mathsf{R}_{M_{j},k}$ instead of $\mathsf{R}_{k}$ for the sake of clarity.
@@
We recall that if a pointed non-compact manifold $(M,g,p)$ has non-negative Ricci curvature, then Gromov pre-compactness theorem ensures that, for every sequence $(M,g/r_{j},p)$ as $r_{j}\to+\infty$, there exists a subsequence converging in pGH sense to a metric space $(X,\mathsf{d},p_{\infty})$. Such spaces obtained via blow-downs are called tangent cones at infinity of $M$.
Remark 3.9 (Manifolds with asymptotically non-negative Ricci curvature).
Let $(M,g,p)$ be a manifold with asymptotically non-negative Ricci curvature, i.e.
$\liminf_{x\to\infty}\mathsf{d}(p,x)^{2}\mathsf{Ric}(x)\wedge 0=0.$ (17)
Let $r_{i}\uparrow+\infty$ be a sequence of scales. Let $c>0$. In the rescaled spaces $(M_{i},g/r_{i}^{2},p)$ (where $M=M_{i}$ as sets), consider $x\notin B^{i}_{c}(p)\subset M_{i}$. Then, using $\mathsf{d}(p,x)\geq cr_{i}$, it holds
$\mathsf{Ric}_{M_{i}}(x)\wedge 0=r_{i}^{2}\mathsf{Ric}_{M}(x)\wedge 0\geq c^{-2}\mathsf{d}(p,x)^{2}\mathsf{Ric}_{M}(x)\wedge 0.$
It follows from (17) that, for every $\delta>0$, the rescaled spaces $(M_{i},g/r_{i}^{2},p)$ have $\mathsf{Ric}_{M_{i}}\geq-\delta$ outside of $B^{i}_{c}(p)$ for $i$ large enough.
In particular, for every $c>0$, the spaces $M_{i}$ are all uniformly locally doubling outside of $B^{i}_{c}(p)$. Hence, by Gromov’s precompactness Theorem, up to passing to a subsequence, there exists a pGH limit $(X_{\infty},\mathsf{d}_{\infty},x_{\infty})$. Hence, assumption (17) is enough to guarantee the existence of tangent cones at infinity.
Lemma 3.10.
Let $(M^{n},g,p)$ be a Riemannian manifold with asymptotically non-negative Ricci curvature, i.e.
$\liminf_{x\to\infty}\mathsf{d}(p,x)^{2}\mathsf{Ric}(x)\wedge 0=0.$
Let $s\in(0,1)$. There exists $c(s,n)>0$ such that
$\limsup_{r\to+\infty}\frac{\mathsf{Vol}(B_{r}(p))}{\mathsf{Vol}(B_{rs}(p))}<c(s,n).$
Proof.
Let $r_{i}\uparrow+\infty$ and consider the rescaled spaces $(M_{i},g/r_{i}^{2},p)$, where $M_{i}=M$ as sets. The statement of the lemma is equivalent to proving
$\limsup_{i\to+\infty}\frac{\mathsf{Vol}_{i}(B^{i}_{1}(p))}{\mathsf{Vol}_{i}(B^{i}_{s}(p))}<c(s,n).$
By Remark 3.9, for every $\delta>0$, for $i$ large enough, it holds
$\mathsf{Ric}_{M_{i}}\geq-\delta\quad\text{on }M_{i}\setminus B^{i}_{s/2}(p).$ (18)
Consider now a collection of balls $\{B^{i}_{(9/100)s}(a_{i,j})\}_{j}$in $M_{i}$ such that
@@
    $a_{i,j}\in B^{i}_{(101/100)s}(p)\setminus B^{i}_{s}(p)$.
@@
    $\{B^{i}_{(9/100)s}(a_{i,j})\}_{j}$ cover $B^{i}_{(101/100)s}(p)\setminus B^{i}_{s}(p)$.
  3. 3.
    $\{B^{i}_{(3/100)s}(a_{i,j})\}_{j}$ are disjoint.
By item 1, there exist points
$b_{i,j}\in B^{i}_{(2/100)s}(a_{i})\cap[B^{i}_{(99/100)s}(p)\setminus B^{i}_{s/2}(p)].$
Since $B^{i}_{(9/100)s}(a_{i,j})\subset B^{i}_{(11/100)s}(b_{i,j})$, by item 2, the collection $\{B^{i}_{(11/100)s}(b_{i,j})\}_{j}$ covers $B^{i}_{(101/100)s}(p)\setminus B^{i}_{s}(p)$. Hence,
$\mathsf{Vol}_{i}(B^{i}_{(101/100)s}(p)\setminus B^{i}_{s}(p))\leq\sum_{j}\mathsf{Vol}_{i}(B^{i}_{(11/100)s}(b_{i,j})).$ (19)
Since $B^{i}_{(11/100)s}(b_{i,j})\subset M_{i}\setminus B^{i}_{s/2}(p)$, by (18) and Bishop-Gromov’s inequality, there exists a constant $c(n)$ such that, for $i$ large enough,
$\sum_{j}\mathsf{Vol}_{i}(B^{i}_{(11/100)s}(b_{i,j}))\leq c(n)\sum_{j}\mathsf{Vol}_{i}(B^{i}_{(1/100)s}(b_{i,j})).$ (20)
Since $B^{i}_{(1/100)s}(b_{i,j})\subset B^{i}_{(3/100)s}(a_{i,j})$, by item 3, the collection $\{B^{i}_{(1/100)s}(b_{i,j})\}_{j}$ consists of disjoint balls. Moreover, by our choice of $b_{i,j}$, it holds $B^{i}_{(1/100)s}(b_{i,j})\subset B^{i}_{s}(p)$. Hence,
$\sum_{j}\mathsf{Vol}_{i}(B^{i}_{(1/100)s}(b_{i,j}))\leq\mathsf{Vol}_{i}(B^{i}_{s}(p)).$ (21)
Combining (19), (20), and (21), for $i$ large enough, it holds
$\mathsf{Vol}_{i}(B^{i}_{(101/100)s}(p))\leq c(n)\mathsf{Vol}_{i}(B^{i}_{s}(p)).$
Iterating this procedure a sufficient number of times, we find
$\frac{\mathsf{Vol}_{i}(B^{i}_{1}(p))}{\mathsf{Vol}_{i}(B^{i}_{s}(p))}\leq c(n,s),$
concluding the proof. ∎
In order to work on manifolds with asymptotically non-negative Ricci curvature, we need slightly modified versions of Theorems 1.1 and 3.8 (i.e., Theorems 3.12 and 3.13). These will follow repeating the previous proofs, once Theorem 3.11 below is established.
Theorem 3.11 is a modified version of item 2 in Theorem 2.6, where the lower bound on the Ricci curvature is now only local, but the closedness to Euclidean space is required at large scales. Another difference is that the maps provided by Theorem 3.11 are not strictly speaking $(k,\delta)$-splitting, since their gradient is bounded just by a dimensional constant which, a priori, may not be close to $1$. Nevertheless, these maps are regular enough to prove Theorem 3.12.
Let us remark that Theorem 3.11 does not follow by repeating the proof of Theorem 2.6 (see [CJN, Section 4.9]), as the latter uses the lower semicontinuity of Cheeger energies under pmGH convergence of spaces, which relies on global Ricci curvature lower bounds. Instead, Theorem 3.11 can be proved along the lines of the Almost-Splitting Theorem from [ChCo1]. We only sketch the proof.
Theorem 3.11.
Let $k\leq n$ be natural numbers. For every $\epsilon>0$, there exists $\delta>0$ such that the following hold. If $(M^{n},g,p)$ is a manifold with $\mathsf{Ric}_{M}\geq-\delta$ on $B_{\delta^{-1}}(p)$ and $\mathsf{d}_{GH}(B^{M}_{\delta^{-1}}(p),B^{\mathbb{R}^{k}}_{\delta^{-1}}(0))<\delta$, then there exists an harmonic map $\phi:B_{1}(p)\to\mathbb{R}^{k}$ with
  1. 1.
    $\fint_{B_{1}(p)}|\nabla\phi_{i}\cdot\nabla\phi_{j}-\delta_{ij}|\,{\rm d}\mathsf{Vol}\leq\epsilon$ for every $i,j$.
  2. 2.
    $|\nabla\phi|\leq c(n)$.
  3. 3.
    $\fint_{B_{1}(p)}|\mathsf{Hess}(\phi)|^{2}\,{\rm d}\mathsf{Vol}\leq\epsilon$.
Proof.
Fix $\epsilon>0$ and let $0<\delta\ll 1$ to be chosen later. Let $\{e_{i}\}_{i=1}^{k}$ be an orthonormal basis of $\mathbb{R}^{k}$. For every $i\in\{1,\cdots,k\}$, let $p_{i}\in B^{M}_{\delta^{-1}}(p)$ be a point whose image via the $\delta$-GH map from $B^{M}_{\delta^{-1}}(p)$ to $B^{\mathbb{R}^{k}}_{\delta^{-1}}(0)$ is $\delta$-close to $\frac{\delta^{-1}}{2}e_{i}$. Consider the map $b_{i}:M\to\mathbb{R}$ defined by
$b_{i}(x):=\mathsf{d}(x,p_{i})-\mathsf{d}(p,p_{i}).$
Observe that for every $R,\epsilon^{\prime}>0$, it holds $\Delta b_{i}\leq\epsilon^{\prime}$ in $B^{M}_{R}(p)$ if $\delta$ is small enough. Consider then the harmonic replacement $h_{i}\in C^{\infty}(B^{M}_{100}(p))$ (assuming $\partial B^{M}_{100}(p)$ to be smooth, or approximating it with smooth sets otherwise) given by
$\Delta h_{i}=0\quad\text{on }B^{M}_{100}(p),\quad h_{i}=b_{i}\quad\text{on }\partial B^{M}_{100}(p).$
We claim that the restriction of the map $\phi:=(h_{1},\cdots,h_{k}):B^{M}_{100}(p)\to\mathbb{R}^{k}$ to $B^{M}_{1}(p)$ has the desired properties. By [CheegerLectures, Lemma 9.13] (after [ChCo1]), if $\delta>0$ is small enough, we have
$\fint_{B_{1}(p)}|\mathsf{Hess}(\phi)|^{2}\,{\rm d}\mathsf{Vol}\leq\epsilon.$
By [CheegerLectures, Lemma 9.10] (again after [ChCo1]) and [ColdingConv, Lemma 2.9] (which holds even if $k<n$), if $\delta>0$ is small enough, we have
$\fint_{B_{1}(p)}|\nabla h_{i}\cdot\nabla h_{j}-\delta_{ij}|\,{\rm d}\mathsf{Vol}\leq\epsilon.$
The bound on the gradient of $\phi$ follows by Cheng-Yau’s gradient estimates [ChengYau]. ∎
The proof of the next result is the same as the one of Theorem 1.1, replacing the use of Theorem 2.6 with Theorem 3.11.
Theorem 3.12.
Let $L,k,n\in\mathbb{N}$ with $k\leq n$ be fixed. Then,
$\sup\Big{\{}\fint_{B_{1}(x)}\mathsf{R}_{k}\wedge L\,{\rm d}\mathsf{Vol}:(M,g,x)\text{ has }\mathsf{Ric}_{M}\geq-\delta\text{ on }B_{\delta^{-1}}(x),\text{ and }\mathsf{d}_{pGH}(M,\mathbb{R}^{k})<\delta\Big{\}}$
converges to 00 as $\delta\searrow 0$.
By repeating the proof of item 2 in Theorem 3.8 (and using Theorem 3.12 instead of Theorem 1.1), one obtains the following result.
Theorem 3.13.
Let $\epsilon,L\in(0,+\infty)$ and let $n\in\mathbb{N}$. There exists $\delta>0$, such that the following holds. Let $(M^{n},g,p)$ be a manifold with $\mathsf{Ric}_{M}\geq-\delta$ on $B_{\delta^{-1}}(p)$, and
Then, $\mathsf{d}_{\rm{pGH}}(M,\mathbb{R}^{n-1})\geq\delta$ and $\mathsf{d}_{\rm{pGH}}(M,\mathbb{R}^{n})\geq\delta$.
We now prove Theorem 1.4 from the Introduction.
Theorem 3.14.
Let $(M^{n},g,p)$ be a Riemannian manifold and let $L>0$ and $\alpha\in(0,2)$.
  1. 1.
    Assume that
    $\liminf_{x\to\infty}\mathsf{d}(p,x)^{2}\mathsf{Ric}(x)\wedge 0=0\quad\text{and}\quad\lim_{r\to+\infty}\fint_{B_{r}(p)}0\vee(r^{2-\alpha}\mathsf{R}_{k})\wedge L\,{\rm d}\mathsf{Vol}=L.$
@@
    Let $(X,\mathsf{d},x_{\infty})$ be any tangent cone at infinity of $M$. If $(\mathbb{R}^{d},\mathsf{d}_{eu})$ is a tangent space to $X$ at a point $x\neq x_{\infty}$, then $d\leq k-1$.
  2. 2.
    Assume that
    $\liminf_{x\to\infty}\mathsf{d}(p,x)^{2}\mathsf{Ric}_{n-2}(x)\wedge 0=0\quad\text{and}\quad\lim_{r\to+\infty}\fint_{B_{r}(p)}0\vee(r^{2-\alpha}\mathsf{R})\wedge L\,{\rm d}\mathsf{Vol}=L.$
    Let $(X,\mathsf{d},x_{\infty})$ be any tangent cone at infinity of $M$. If $(\mathbb{R}^{d},\mathsf{d}_{eu})$ is a tangent space to $X$ at a point $x\neq x_{\infty}$, then $d\leq n-2$.

### v3

@@ 2 Preliminaries
The next proposition, from [37, Theorem $1.1$], characterizes the $\mathsf{RCD}$ spaces having essential dimension $1$.
@@
We now recall some topological properties of $\mathsf{RCD}$ spaces. The next proposition is a special case of [49, Theorem 5], in the non-collapsing scenario. We denote by $\mathrm{b}_{1}(X)$ the first Betti number of $X$.
@@
The next proposition, proved in [57, Corollary 1.2] (see also [58, Theorem 1.2]), establishes the converse inequality in higher generality.
@@
[…]
The next result proved in [43] establishes the existence of the universal cover for an $\mathsf{RCD}(K,N)$ space.
@@
[…]
Since $\mathsf{RCD}(K,N)$ spaces are semi-locally simply connected thanks to [57] the universal is simply connected. The next proposition follows from [43, Theorem 1.3], after taking into account that the revised fundamental group mentioned in the reference coincides with the usual fundamental group thanks to [57].
@@
The next proposition from [51, Corollary 4.11] shows that the universal cover of a Ricci limit space is a Ricci limit space, as well.
@@
[…]
The next proposition follows combining [3, Theorems $1.1$ and $1.3$] with the fact that each covering of a fixed manifold has volume growth controlled from above by the one of the universal cover.
@@
[…]
@@ 3 Integral estimates on scalar curvature and first geometric applications
Let $k,n\in\mathbb{N}$ with $k\leq n$, and let $c,\epsilon>0$ be fixed. There exist $c_{1}(n,\epsilon),\epsilon^{\prime}(n,\epsilon)>0$ arbitrarily small satisfying the following. Let $A=(a_{j,i})\in\mathbb{R}^{n\times k}$ be a matrix with $|A^{T}A-I_{k}|\leq c_{1}$ and $|A|^{2}\leq c$. Consider real numbers $\epsilon^{\prime}\leq\lambda_{1}\leq\cdots\leq\lambda_{n}$ with $\lambda_{k}\geq\epsilon/2k$. Then
@@
Assume by contradiction that the statement fails. Hence, we find sequences $c_{h},\epsilon^{\prime}_{h}\downarrow 0$ and $A_{h}=(a^{h}_{j,i})\in\mathbb{R}^{n\times k}$ with $|A^{T}_{h}A_{h}-I_{k}|\leq c_{h}$, $|A_{h}|^{2}\leq c$ and real numbers $\epsilon^{\prime}_{h}\leq\lambda^{h}_{1}\leq\cdots\leq\lambda^{h}_{n}$ with $\lambda_{k}\geq\epsilon/2k$, with
@@
Moreover, since $|A_{h}|^{2}\leq c$, $\epsilon^{\prime}_{h}\leq\lambda^{h}_{1}\leq\cdots\leq\lambda^{h}_{n}$ and $\lambda^{h}_{k}\geq\epsilon/2k$, it holds
@@
The last two identities are clearly in contradiction: indeed, the former states that the columns of $A$ are $k$ orthonormal vectors in $\mathbb{R}^{n}$, while the latter implies that their span is at most k1k-1 dimensional. ∎
@@
Observe that the condition in the previous definition is independent of the chosen orthonormal basis for $T_{p}M$. Fix now an orthonormal basis in $T_{p}M$ such that $\mathsf{Ric}_{M}$ in $T_{p}M$ is represented by the diagonal matrix $\Lambda\in\mathbb{R}^{n\times n}$, whose eigenvalues are $\epsilon^{\prime}\leq\lambda_{1}\leq\cdots\leq\lambda_{n}$.
@@
$0$ $\vee\sum_{j=1}^{k}\mathsf{Ric}(\nabla u_{j},\nabla u_{j})=0\vee\sum_{j=1}^{k}\nabla u_{j}^{T}\Lambda\nabla u_{j}$
@@
We first prove that $E^{\prime}\subset E$. It is sufficient to show that at every point of $E^{\prime}$ the quantity $|\nabla u^{T}\nabla u-I_{k}|$ is smaller than the constant $c_{1}(\epsilon,n)$ appearing in the definition of $E$. We prove it by a telescopic argument (cf. [5]). Set $\mathcal{E}:=\sum_{i,j}|\nabla u_{j}\cdot\nabla u_{j}-\delta_{ij}|$. For every $j\in\mathbb{N}$ and $y\in B_{1}(x)$, we have
@@
Choosing $\eta$ and $\epsilon^{\prime}$ to be sufficiently small, and letting $j$ go to $\infty$, we conclude that $E^{\prime}\subset E$.
@@
To prove this, we use a Vitali covering argument coupled with a scale-picking (cf. [5]). For every $y\in B_{1}(x)$ such that
@@
$\sup\Big\{\fint_{B_{1}(x)}\mathsf{R}_{k}\wedge L\,{\rm d}\mathsf{Vol}:(M,g,x)\text{ has }\mathsf{Ric}_{M}\geq-\delta,\text{ and }B_{10}(x)\text{ is $(k,\delta)$-symmetric}\Big\},$ (10)
$\sup\Big\{\fint_{B_{1}(x)}|\mathsf{R}|^{s}\,{\rm d}\mathsf{Vol}:(M,g,x)\text{ has }\mathsf{Ric}_{M}\geq-\delta,\text{ and }B_{10}(x)\text{ is }\delta\text{-regular}\Big\}$ (11)
converge to 0 as $\delta\searrow 0$.
@@
Let $k,n\in\mathbb{N}$ with $k<n$. We show that for every $0<\epsilon<1/2$ there exists $\delta>0$ such that if $(M^{n},g)$ has $\mathsf{Ric}_{M}\geq-\delta$ and $B_{10}(x)\subset M$ is $(k,\delta)$-symmetric, then
@@
We next show (11). To this aim, fix $s\in(0,1)$ and consider the case $k=n$. We want show that for every $0<\epsilon<1/2$ there exists $\delta>0$ such that if $(M^{n},g,x)$ has $\mathsf{Ric}_{M}\geq-\delta$ and $B_{10}(x)\subset M$ is $\delta$-regular, then
@@
$\sup\Big\{\fint_{B_{1}(x)}|\mathsf{R}|\,{\rm d}\mathsf{Vol}:(M,g,x)\text{ has }\mathsf{Ric}_{M}\geq-\delta\text{ and }B_{10}(x)\text{ is }\delta\text{-regular }\Big\}\to 0$
as $\delta\to 0$. This would allow to obtain a priori integral bounds on the scalar curvature outside of the quantitative singular set of a manifold with Ricci curvature bounded from below. Unfortunately, it is shown in [21] that such estimates cannot exist in such generality.
@@ 3.2 First geometric applications in almost non-negative Ricci and integrally-positive scalar curvature
We now obtain the first geometric applications of the results of the previous section. We introduce some notation first (cf. [50]).
@@
Let $(M^{n},g)$ be a manifold, let $1\leq k\leq n-1$, let $x\in M$, and let $c\in\mathbb{R}$. We denote by $R_{M}(\cdot,\cdot,\cdot,\cdot)$ the Riemann curvature tensor of $M$. We say that $\mathsf{Ric}_{M,k}\geq c$ (resp. $\leq c$) in $T_{x}M$ if, for all $(k+1)$-dimensional subspaces $V\subset T_{x}M$, it holds
@@
Let $(M^{n},g)$ be a manifold, let $1\leq k\leq n-1$, let $A\subset M$ be measurable, let $c\in\mathbb{R}$, and let $F:\mathbb{R}\to\mathbb{R}$ be a measurable function. We say that
$\int_{A}F(\mathsf{Ric}_{M,k})\,d\mathsf{Vol}\geq c$
@@
We say that $\int_{A}F(\mathsf{Sec}_{M})\,d\mathsf{Vol}\geq c$ if $\int_{A}F(\mathsf{Ric}_{1,M})\,d\mathsf{Vol}\geq c$.
The next theorem should be compared with [59, Theorem 1.1], where similar conclusions were obtained under stronger assumptions: non-negative Ricci curvature (below, almost non-negative Ricci curvature) and scalar curvature greater than 2 (below, a local integral lower bound on the scalar curvature). Whenever we consider a sequence of manifolds $(M_{j},g_{j})$, we write $\mathsf{R}_{M_{j},k}$ instead of $\mathsf{R}_{k}$ for the sake of clarity.
@@
We recall that if a pointed non-compact manifold $(M,g,p)$ has non-negative Ricci curvature, then Gromov pre-compactness theorem ensures that, for every sequence $(M,g/r_{j},p)$ as $r_{j}\to+\infty$, there exists a subsequence converging in pGH sense to a metric space $(X,\mathsf{d},p_{\infty})$. Such spaces obtained via blow-downs are called tangent cones at infinity of $M$. We now prove Theorem 1.4 from the Introduction.
Theorem 3.9.
Let $(M^{n},g,p)$ be a Riemannian manifold; let $L>0$ and $\alpha\in(0,2)$.
@@
    Assume that $\mathsf{Ric}\geq 0$ and
    $\lim_{r\to+\infty}\fint_{B_{r}(p)}0\vee(r^{2-\alpha}\mathsf{R}_{k})\wedge L\,{\rm d}\mathsf{Vol}=L.$
    Let $(X,\mathsf{d},x_{\infty})$ be any tangent cone at infinity of $M$. If $(\mathbb{R}^{d},\mathsf{d}_{eu})$ is a tangent space to $X$ at a point $x\in X$, then $d\leq k-1$.
@@
    Assume that $\mathsf{Ric}_{n-2}\geq 0$ and
    $\lim_{r\to+\infty}\fint_{B_{r}(p)}0\vee(r^{2-\alpha}\mathsf{R})\wedge L\,{\rm d}\mathsf{Vol}=L.$
@@
    Let $(X,\mathsf{d},x_{\infty})$ be any tangent cone at infinity of $M$. If $(\mathbb{R}^{d},\mathsf{d}_{eu})$ is a tangent space to $X$ at a point $x\in X$, then $d\leq n-2$.